Pam and Sue are both driving. After 4 hours, they are 440 miles apart. Pam is driving 10mph faster than Sue, how fast is Sue driving?
Found 3 solutions by rfer, TimothyLamb, Alan3354:
Answer by rfer(16322)

440/4=110
x+x+10=110
2x=110-10
x=100/2
x=50
x+10=60

Answer by TimothyLamb(4379)

s = d / t
---
p + s = 440/4
p + s = 110
---
p = s + 10
---
p + s = 110
s + 10 + s = 110
2s = 100
---
s = 50 mph
p = 60 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by Alan3354(69443)

Depends on the direction they're driving, and their starting points.